OpenStudy (anonymous):

Solve: 3(x + 4) - 4(x + 3) = 7x + 16. @Pre-Algebra

OpenStudy (aroub):

Distribute: 3x+12-4x-12=7x+16 solve for x: => -7x-4x+3x=-12+12+16 => -8x=16 => x=-2 :D

OpenStudy (anonymous):

x=-2